OpenStudy (anonymous):

Can anyone explain to me why 6^sqrt(2)*6^sqrt(2) = 36^sqrt(2) and not 6^2sqrt(2)?

OpenStudy (anonymous):

You know, since b^n * b^m = b^n+m

OpenStudy (anonymous):

Hello, KarenKaren!

OpenStudy (anonymous):

36^sqrt(2) = 6^(2sqrt(2)) your answer is valid

OpenStudy (anonymous):

Ohhh...

OpenStudy (anonymous):

But wait...what distinguishes the two?